### v1

@@ Appendix B The initial point $x^{0}$ given by (72) satisfies the condition (41)
Firstly, we shall show $\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}\neq 0$, ensuring that the $\hat{\theta}$ given by (71) is well-defined. Note that $(b-Tb)^{\top}A\neq 0$, since $A$ has full row-rank and $b\neq Tb$ follows from the fact that $b\notin\mathcal{S}_{\mu}$. Then we derive $|(b-Tb)^{\top}a_{i_{\star}}|>0$ immediately, which along with the fact $b^{\top}(a_{i_{\star}}-Ta_{i_{\star}})=(b-Tb)^{\top}a_{i_{\star}}$ implies that $a_{i_{\star}}-Ta_{i_{\star}}\neq 0$. Hence, we deduce that $\|a_{i_{\star}}\|^{2}_{2}-\|Ta_{i_{\star}}\|^{2}_{2}>0$ and $\hat{\theta}\neq 0$. Secondly, due to $\underline{x}<0<\overline{x}$ and $\hat{\theta}\neq 0$, the $x^{0}$ given by (72) is feasible for problems (2) and (3), that is, $x^{0}\neq 0$ and $\underline{x}\leq x^{0}\leq\overline{x}$ hold. Thirdly, owing to $0\in[\underline{x}_{i_{\star}},\overline{x}_{i_{\star}}]$ and $x_{i_{\star}}^{0}\neq 0$, we deduce from (72) that $(x_{i_{\star}}^{0}-\hat{\theta})^{2}<\hat{\theta}^{2}$, or equally, $(x_{i_{\star}}^{0}-2\hat{\theta})x_{i_{\star}}^{0}<0$. Note that $\|\mathcal{T}_{\mu}(Ax-b)\|^{2}_{2}=\max\left\{\|E(Ax-b)\|^{2}_{2}:E\in\{0,1\}^{m\times m}\text{ is diagonal,}~{}\|E\|_{0}\leq\mu\right\},$ where $\|\cdot\|_{0}$ counting the number of nonzero elements in the matrix. Then, we have
@@
<$\overset{(*)}{<}1+\frac{\lambda}{2}(\|b\|_{2}^{2}-\|Tb\|_{2}^{2})=1+\frac{\lambda}{2}\mathrm{dist}^{2}(b,\mathcal{S}_{\mu}),$
where ()(*)( ∗ ) follows from the positivity of $\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}$ and the negativity of $(x_{i_{\star}}^{0}-2\hat{\theta})x_{i_{\star}}^{0}$. This means that $x^{0}$, defined in (72), satisfies the condition (41).

### v2

@@ Appendix B The initial point $x^{0}$ given by (77) satisfies the condition (43)
Firstly, we shall show $\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}\neq 0$, ensuring that the $\hat{\theta}$ given by (76) is well-defined. Note that $(b-Tb)^{\top}A\neq 0$, since $A$ has full row-rank and $b\neq Tb$ follows from the fact that $b\notin\mathcal{S}_{\mu}$. Then, according to the definition of $i_{\star}$, we have $|(b-Tb)^{\top}a_{i_{\star}}|>0$, which along with the fact $b^{\top}(a_{i_{\star}}-Ta_{i_{\star}})=(b-Tb)^{\top}a_{i_{\star}}$ implies that $a_{i_{\star}}-Ta_{i_{\star}}\neq 0$. Hence, we deduce that $\|a_{i_{\star}}\|^{2}_{2}-\|Ta_{i_{\star}}\|^{2}_{2}>0$ and $\hat{\theta}\neq 0$. Secondly, due to $\underline{x}<0<\overline{x}$ and $\hat{\theta}\neq 0$, the $x^{0}$ given by (77) is feasible for problems (2) and (3), that is, $x^{0}\neq 0$ and $\underline{x}\leq x^{0}\leq\overline{x}$ hold. Thirdly, owing to $0\in[\underline{x}_{i_{\star}},\overline{x}_{i_{\star}}]$ and $x_{i_{\star}}^{0}\neq 0$, we deduce from (77) that $(x_{i_{\star}}^{0}-\hat{\theta})^{2}<\hat{\theta}^{2}$, or equally, $(x_{i_{\star}}^{0}-2\hat{\theta})x_{i_{\star}}^{0}<0$. Note that $\|\mathcal{T}_{\mu}(Ax-b)\|^{2}_{2}=\max\left\{\|E(Ax-b)\|^{2}_{2}:E\in\{0,1\}^{m\times m}\text{ is diagonal,}~\|E\|_{0}\leq\mu\right\},$ where $\|\cdot\|_{0}$ counting the number of nonzero elements in the matrix. Then, we have
@@
$\overset{(*)}{<}1+\frac{\lambda}{2}(\|b\|_{2}^{2}-\|Tb\|_{2}^{2})=1+\frac{\lambda}{2}\mathrm{dist}^{2}(b,\mathcal{S}_{\mu}),$
where ()(*) follows from the positivity of $\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}$ and the negativity of $(x_{i_{\star}}^{0}-2\hat{\theta})x_{i_{\star}}^{0}$. This means that $x^{0}$, defined in (77), satisfies the condition (43).
Funding   The work of Na Zhang was supported in part by the National Natural Science Foundation of China under grant 12271181, by the Guangzhou Basic Research Program under grant 2025A04J5240 and by the Basic and Applied Basic Research Foundation of Guangdong Province under grant 2023A1515030046.
The work of Qia Li was supported in part by the National Natural Science Foundation of China under grant 12471098 and the Guangdong Province Key Laboratory of Computational Science at the Sun Yat-sen University (2020B1212060032).
Data Availability   The datasets generated or analyzed during the current study are available from the corresponding author on reasonable request.
Declarations
Conflict of Interest   The authors declare that they have no conflict of interest.